ten times the sum of twelve and a number is 27. whats the number
Answer by ashleyk_xox(2)

ten times the sum of twelve and a number is 27. whats the number)
10(x+12)=27 The sum of twelve and a number (x+12) times 10
10x+120=27 Move 120 to the right side which makes it -120
(27-120=-93)
10x=-93
x=-93/10
x=-9.3
Check your answer
(12+(-9.3)10=27
12-9.3=2.7
2.7*10=27
27=27
hope that helps!